# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 3 Applications
In the classic submodular cover problem [Wol82], there is a monotone submodular function $f:2^{E}\rightarrow\mathbb{Z}_{+}$ defined on a groundset $E$, where each item $i\in E$ has some cost $c_{i}$. We assume that the function is integer-valued with maximal value $Q$. The goal is to select a subset $S\subseteq E$ that “covers” the function, i.e., $f(S)=Q$. The objective is to minimize the total cost $\sum_{i\in S}c_{i}$ of the chosen items. Note that the submodular cover problem generalizes set cover. It is well-known that this problem can be approximated within a factor $1+\ln(Q)$ [Wol82], and no better approximation is possible [DS14]. We consider a stochastic version, where each item $i\in E$ is active independently with probability $p_{i}$, and the function value is accrued only on active items. The goal now is to select items sequentially (and adaptively) so as to cover function $f$ (or run out of items to select), at the minimum expected cost. This problem has been studied extensively [GK17, INZ16, HKP21], and a $1+\ln(Q)$ approximation algorithm is also known for it [HKP21]; this result even holds in the more general formulation with multiple outcomes for each item. Using the fact that the function $f$ is monotone, it is easy to see that stochastic submodular cover is up-monotone. So, using Theorem 1.1 with $k=2$, we obtain a polynomial time online learning algorithm for stochastic submodular cover having $(1+\ln Q)$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
@@ Appendix A Sampling-based Algorithms
A stronger monotonicity condition is that of strong monotonicity [GHTZ21], defined next.
@@
A stochastic problem is strongly up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{{E}}$, we have $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{D}}},\mathbf{{x}})]\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$, where $\sigma_{\mathbf{{D}}}$ is the optimal policy for instance $\mathcal{I}$ under distribution $\mathbf{{D}}$. The problem is said to be strongly down-monotone if $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{D}}},\mathbf{{x}})]\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ under the same conditions as above.
We note that strong monotonicity implies monotonicity. In prior work, [GHTZ21] gave optimal sample complexity bounds for stochastic optimization problems that exhibit the strong monotonicity condition. They also proved this property for problems including Prophet inequality and Pandora’s box. We make use of this property in §3, when we apply our result to these problems.
Below, we assume that $f_{max}=1$ by scaling. The sample-complexity result in [GHTZ21] is:
Theorem A.1 (Theorem 17 [GHTZ21]).
@@
Recall that, in the full-feedback model, we get one sample from each r.v. in every time-step. We now describe the algorithm. The algorithm is straightforward: for each $t=1,2,\ldots T$, we use estimates from the prior $t-1$ time steps to obtain a policy to use for the $t^{\text{th}}$ time-step. See Algorithm 4 for a formal description.
Algorithm 4 Exploit
1:Input: one sample $\mathbf{{x}}=(x_{1},\cdots,x_{n})\sim\mathbf{{D}}$,
2:for $t=1,2,\ldots,T$ do
3:     let $\mathbf{{\widehat{D}_{t}}}\leftarrow$ frequentist estimate from previous $t-1$ time-steps
4:     run policy $\mathtt{OPT}(\mathbf{{\widehat{D}_{t}}})$ for the $t^{\text{th}}$ time-step
@@
$\mathbb{E}[R_{T}]\leq\mathbb{E}[R_{T}\mid G]\cdot\mathbf{P}(G)+\mathbb{E}[R_{T}\mid\overline{G}]\cdot\mathbf{P}(\overline{G})\leq O\left(\sqrt{nT\log(nT)}\right)+T\cdot\frac{1}{T}=O\left(\sqrt{nT\log(nT)}\right).$
@@
A significant challenge in the semi-bandit feedback model arises from our lack of control over the r.v.s from which we get samples. One potential strategy to address this issue is to artificially generate samples for an item $i$ by probing item $i$ first in the algorithm’s policy for that period. However, this approach comes with an inherent drawback – probing item $i$ first may result in a poor policy, and so we suffer a high regret in such periods. The standard explore-then-exploit algorithm that first gets $T^{2/3}$ samples for each r.v., and then plays the optimal policy (for the empirical distribution) for the remaining time steps only achieves regret $\widetilde{O}(T^{2/3})$.
In the hope of obtaining improved regret guarantees, we consider a different algorithm where we pair periods of exploration with some periods of exploitation, wherein we leverage the learned distribution. To do this, we divide the time horizon $T$ into meta periods, each of length $n+\beta n$ (we optimize $\beta$ later). In meta-period $h$, we first gather $n$ samples (one for each r.v.) and then use policy $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$ for $\beta n$ time-steps. Here $\mathbf{{\widehat{D}_{h}}}$ denotes the frequentist distribution obtained from the $h$ samples (until meta-period $h$) of $\mathbf{{D}}$. Note that we have no control over the r.v.s probed by $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$, and thus cannot use these samples in our analysis.
By Theorem A.1, with probability at least $1-\frac{1}{T^{2}}$, we have that $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$ is an $\epsilon_{h}$-additive approximation to $\mathtt{OPT}(\mathbf{{D}})$ where $\epsilon_{h}\leq\sqrt{\frac{Cn\log(nT)}{h}}$. By union bound, this is true for all meta-periods $h=0,1,\ldots,$ with probability at least $1-\frac{1}{T}$ (since $h\leq T$). As before, we consider this to be a good event, $G$. Under this event, the regret in meta-period $h$ is at most $n+\beta n\cdot\sqrt{\frac{Cn\log(nT)}{h}}$, and the total regret under $G$ is at most
$nH+C\beta n^{3/2}\sqrt{\log(nT)}\sum_{h=1}^{H}\sqrt{\frac{1}{h}}=nH+C\beta n^{3/2}\sqrt{H\log(nT)}$
where $H=\frac{T}{(1+\beta)n}$. The optimal choice for $\beta=O\left(\left(\frac{T}{n^{2}\log T}\right)^{1/3}\right)$, and so we again end up with a regret of $O\left((nT)^{2/3}(\log T)^{1/3}\right)$. To summarize, we are not aware of any generic approach that reduces sample-complexity bounds to regret minimization in the semi-bandit have $o(T^{2/3})$ regret.
@@
We note that [GHTZ21] also give (slightly worse) sample complexity bounds for a broader class of problems that need not satisfy any monotonicity property, but have a finite support-size $k$. Using the approaches described above, we can convert these sample-complexity guarantees to obtain $O\left(\sqrt{nkT\log(nT)}\right)$ regret under full-feedback and $O\left((nkT)^{2/3}(\log T)^{1/3}\right)$ regret under semi-bandit (and bandit) feedback.

### v2

@@ 5.7 Stochastic Submodular Optimization
In the classic submodular cover problem (Wolsey, 1982), there is a monotone submodular function $f:2^{E}\rightarrow\mathbb{Z}_{+}$ defined on a groundset $E$, where each item $i\in E$ has some cost $c_{i}$. We assume that the function is integer-valued with maximal value $Q$. The goal is to select a subset $S\subseteq E$ that “covers” the function, i.e., $f(S)=Q$. The objective is to minimize the total cost $\sum_{i\in S}c_{i}$ of the chosen items. Note that the submodular cover problem generalizes set cover. It is well-known that this problem can be approximated within a factor $1+\ln(Q)$ (Wolsey, 1982), and no better approximation is possible (Dinur and Steurer, 2014). We consider a stochastic version, where each item $i\in E$ is active independently with probability $p_{i}$, and the function value is accrued only on active items. The goal now is to select items sequentially (and adaptively) so as to cover function $f$ (or run out of items to select), at the minimum expected cost. This problem has been studied extensively (Golovin and Krause, 2017; Im et al., 2016; Hellerstein et al., 2021), and Hellerstein et al. (2021) obtained a $1+\ln(Q)$ approximation algorithm for it. Using the fact that the function $f$ is monotone, it is easy to see that stochastic submodular cover is up-monotone. So, using Theorem 1.1 with $k=2$, we obtain a polynomial time online learning algorithm for stochastic submodular cover having $(1+\ln Q)$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
5.8 Single Resource Revenue Management
We now consider the single resource revenue management problem ($\mathsf{SRM}$). A firm (say, airline) has $C$ units of a certain resource (for example, seats on a flight). This resource must be allocated among $n$ fare classes, where each unit sold in class $i$ generates revenue $p_{i}$. Demands for these fare classes arrive sequentially: we use the convention from Gallego et al. (2019) that classes arrive in the order $n,n-1,\ldots,1$. This problem is often studied under the “low before high” assumption that $p_{n}<\cdots<p_{1}$, which reflects the fact that cheaper fares are offered earlier. The demands are random and independent across classes: we use $\mathsf{X}_{i}\sim\mathcal{D}_{i}$ to denote the demand for class $i$. We assume that each $\mathsf{X}_{i}$ is a discrete r.v. with support $\{0,1,\ldots,C\}$. The goal is to decide how much capacity to make available for each fare class so as to maximize the expected total revenue. For each class $i=n,\ldots,1$, the firm needs to first decide on the number $b$ of units offered (which cannot exceed the available units) and then the demand $\mathsf{X}_{i}$ materializes, which reduces the available units by $\min(\mathsf{X}_{i},b)$. When the demand distributions are unknown, censored feedback is the most natural, where we observe only the minimum of the actual demand and the offered capacity. As an application of Theorem 4.1 we obtain a polynomial-time online learning algorithm for $\mathsf{SRM}$ with $1$-regret $O(nC^{2}p_{\max}\sqrt{CT\log(CT)})$ where $p_{\max}$ is the maximum price. This result directly improves the dependence on $T$ compared to Van Ryzin and McGill (2000), whose algorithm achieves a regret of $\widetilde{O}\left(T^{1-\frac{\beta}{2^{n-1}}}\right)$ for some constant $\beta\in(0,1)$. Moreover, it also improves upon the result of Huh and Rusmevichientong (2006), who obtain regret $O(\sqrt{p_{\max}}C^{2n-3}\cdot\sqrt{T})$, thereby removing the exponential dependence on $n$ in the regret bound.
In order to apply Theorem 4.1, we first note that the objective satisfies Assumption 4.1. Indeed, any policy for $\mathsf{SRM}$ involves setting a threshold at each step $i$ corresponding to the number of units offered to class $i$. Moreover, the increase in the $\mathsf{SRM}$ objective at any step is completely determined by the (censored) observation at that step: so the function value at each leaf node of the policy is a constant. Next, we prove that $\mathsf{SRM}$ is down-monotone. This proof relies on some known structural properties of the exact dynamic programming algorithm for $\mathsf{SRM}$ (with known distributions). We now summarize these properties; see Gallego et al. (2019) for more details.
$\mathsf{SRM}$ dynamic program. The value functions in the dynamic program (DP) are as follows. For each $i\in[n]$ and $c\leq C$, let $V_{i}(c)$ denote the optimal total expected revenue from classes $i,i-1,\ldots,1$ given that we have $c$ units of capacity remaining before the arrival of class $i$ demand. The optimal value of the $\mathsf{SRM}$ instance is $\mathtt{OPT}=V_{n}(C)$.
The decision for class $i$ is to determine the number of units offered to class $i$ demand; equivalently, we need to decide the number of units to protect for future classes $i-1,\ldots,1$. Specifically, if $b\in\{0,1,\ldots,c\}$ units are protected, then cbc-b units are available for class $i$ demand. The amount sold to class $i$ equals $\min(c-b,\mathsf{X}_{i})$ and the remaining capacity is $c-\min(c-b,\mathsf{X}_{i})$, which are both random. The revenue generated, also a random variable, is $p_{i}\cdot\min(c-b,\mathsf{X}_{i})$. We let $W_{i}(b,c)$ denote the optimal expected revenue from classes $i,i-1,\ldots,1$ given that we have $c$ units of capacity remaining before the arrival of class $i$ and we protect $b\leq c$ units. Based on the discussion above,
$W_{i}(b,c)=p_{i}\cdot\mathbb{E}[\min(c-b,\mathsf{X}_{i})]+\mathbb{E}[V_{i-1}(c-\min(c-b,\mathsf{X}_{i}))].$ (39)
The first term is the expected revenue obtained from class $i$ given that we offered cbc-b units, and the second term is the expected revenue from classes $i-1,\ldots,1$ when $\max(b,c-\mathsf{X}_{i})$ units of capacity remain. Now, observe that $V_{i}(c)=\max\left\{W_{i}(b,c)\,:\,0\leq b\leq c\right\}$. The base case $V_{1}(c)=p_{1}\cdot\mathbb{E}[\min(c,\mathsf{X}_{1})]$ involves protecting 0 units regardless of $c$ (as there is no demand after this).
We will use the following two properties of this DP (see Proposition 1.5 and Theorem 1.6 in Gallego et al. (2019)).
Lemma 5.5.
Letting $\Delta V_{i}(x):=V_{i}(x)-V_{i}(x-1)$, the value functions satisfy the following.
  • $\Delta V_{i}(x)$ is decreasing in $x=1,\ldots,c$, and
  • $\Delta V_{i}(x)$ is increasing in $i=1,\ldots,n$.
Lemma 5.6.
For each $i\in[n]$ and $c\leq C$, the maximizer of $W_{i}(b,c)$ over $0\leq b\leq c$ equals $\min\{y^{*}_{i-1},c\}$ where
$y^{*}_{i-1}=\max\{0\leq y\leq C:\Delta V_{i-1}(y)>p_{i}\}$
is the protection-level in class $i$. So, the optimal policy is fully specified by $y^{*}_{n-1},y^{*}_{n-2},\ldots,y^{*}_{1},y^{*}_{0}$.
Lemma 5.7.
The single-resource revenue management problem is down-monotone.
Proof.
Consider any instance of the problem with $n$ classes. Let $\mathbf{{D}}=\mathcal{D}_{1}\times\cdots\mathcal{D}_{i}\times\cdots\mathcal{D}_{n}$ and $\mathbf{{E}}=\mathcal{D}_{1}\times\cdots\mathcal{E}_{i}\times\cdots\mathcal{D}_{n}$ be two probability distributions that differ only in the distribution of one class $i$. It suffices to show that $\mathtt{OPT}(\mathbf{{E}})\geq\mathtt{OPT}(\mathbf{{D}})$ whenever $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$. For any $j\in[n]$ and $c\leq C$, let $V_{j}(c)$ and $\overline{V}_{j}(c)$ denote the optimal value functions under distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ respectively. So, $\mathtt{OPT}(\mathbf{{D}})=V_{n}(C)$ and $\mathtt{OPT}(\mathbf{{E}})=\overline{V}_{n}(C)$. In order to prove $V_{n}(C)\leq\overline{V}_{n}(C)$ it suffices to show that $V_{i}(c)\leq\overline{V}_{i}(c)$ for all $c\leq C$, which we do in the rest of the proof.
Let $y_{n-1}^{*},\ldots,y_{1}^{*},y_{0}^{*}=0$ denote the optimal protection levels under product distribution $\mathbf{{D}}$ (from Lemma 5.6). Recall that $V_{i}(c)=W_{i}(\min(y^{*}_{i-1},c),c)$. If $c<y^{*}_{i-1}$ then $V_{i}(c)=V_{i-1}(c)$ and we obtain
$\overline{V}_{i}(c)\geq\overline{V}_{i-1}(c)={V}_{i-1}(c)=V_{i}(c),$
where the first inequality is by ignoring class $i$ and the first equality uses the fact that the demands in classes $i-1,\cdots,1$ have the same distributions in $\mathbf{{D}}$ and $\mathbf{{E}}$.
Below, we assume that $c\geq y^{*}_{i-1}$, which means $V_{i}(c)=W_{i}(y^{*}_{i-1},c)$. To reduce notation, let $y=y^{*}_{i-1}$ and $\Delta_{d}:=\Delta V_{i-1}(d)$ for any $d\leq C$. Let $\widetilde{{\mathsf{X}_{i}}}=\min(c-y,\mathsf{X}_{i})$ be the number of class $i$ units sold, when we implement the optimal policy for $V_{i}(c)$. Note that $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. We have
[…]
$=p_{i}\cdot\mathbb{E}_{\mathcal{D}_{i}}[\widetilde{{\mathsf{X}_{i}}}]+V_{i-1}(y)+\sum_{d=y+1}^{c}\Delta_{d}\cdot\Pr_{\mathcal{D}_{i}}[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
$=p_{i}\cdot(c-y)-p_{i}\cdot\mathbb{E}[c-y-\widetilde{{\mathsf{X}_{i}}}]+V_{i-1}(y)+\sum_{d=y+1}^{c}\Delta_{d}\cdot\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
$=p_{i}\cdot(c-y)+V_{i-1}(y)-p_{i}\sum_{t=1}^{c-y}\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-y-t]+\sum_{d=y+1}^{c}\Delta_{d}\cdot\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
$=p_{i}\cdot(c-y)+V_{i-1}(y)-p_{i}\sum_{d=y+1}^{c}\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]+\sum_{d=y+1}^{c}\Delta_{d}\cdot\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
$=p_{i}\cdot(c-y)+V_{i-1}(y)+\sum_{d=y+1}^{c}\left(\Delta_{d}-p_{i}\right)\cdot\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
Let $\widehat{{\mathsf{X}_{i}}}=\min(c-y,\mathsf{X}_{i})$ where $\mathsf{X}_{i}\sim\mathcal{E}_{i}$ (the class $i$ demand under $\mathbf{{E}}$). We now have
[…]
$=p_{i}\cdot(c-y)+V_{i-1}(y)+\sum_{d=y+1}^{c}\left(\Delta_{d}-p_{i}\right)\cdot\Pr[\widehat{{\mathsf{X}_{i}}}\leq c-d]$
$\geq p_{i}\cdot(c-y)+V_{i-1}(y)+\sum_{d=y+1}^{c}\left(\Delta_{d}-p_{i}\right)\cdot\Pr[\widetilde{{\mathsf{X}_{i}}}\leq c-d]$
The first equality uses $\overline{V}_{i-1}(\cdot)=V_{i-1}(\cdot)$. The second equality follows by the same sequence of steps as for $V_{i}(c)$ above. The last inequality uses the two facts (i) $\widehat{{\mathsf{X}_{i}}}$ stochastically dominates $\widetilde{{\mathsf{X}_{i}}}$ which implies $\Pr[\widehat{{\mathsf{X}_{i}}}\leq z]\leq\Pr[\widetilde{{\mathsf{X}_{i}}}\leq z]$ for all $z\in\mathbb{R}$, and (ii) $\Delta_{d}-p_{i}\leq 0$ for all $d\geq y+1$ by definition of $y^{*}_{i-1}$ in Lemma 5.6. ∎
5.9 Sequential Posted Pricing
We now consider a pricing problem. A seller offers a service to $n$ customers, where each customer’s valuation is modeled as a discrete independent random variable $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. We assume (by scaling) that the values are bounded in $[0,1]$ and that $k$ denotes the maximum support size. The seller must satisfy a feasibility constraint given by a downward-closed set family $\mathcal{F}\subseteq 2^{[n]}$. This means that a subset $S$ of customers can be served if and only if $S\in{\cal F}$. We are interested in designing a truthful mechanism that maximizes the expected revenue. Myerson (1981) provides an optimal mechanism for this class of problems, but it is computationally hard to implement. Therefore, there has been much focus on simpler mechanisms that are approximately optimal. We consider Sequential Posted-Price Mechanisms (SPMs) that were introduced by Chawla et al. (2010). Here, each buyer $i$ is offered a take-it-or-leave-it price $p_{i}$ in some sequence. We consider two variants: Adaptive Sequential Posted-Price Mechanism (ASPM) where the seller can (adaptively) choose the order in which to offer customers, and Fixed Order Sequential Posted-Price Mechanism (FSPM) where the order of customers is fixed and known upfront. Formally, in ASPM, the seller’s policy is represented as a decision tree, where each internal node is labeled with a tuple $(i,p_{i})$, denoting that customer $i$ is offered price $p_{i}$; each node has two outgoing branches corresponding to whether the customer accepts or rejects the offer. In FSPM, the policy only needs to choose the prices $p_{i}$ as the customers will be considered in a fixed order $1,\ldots,n$. The seller collects revenue from all customers who accept their offered price, and the selected set of customers must lie in $\mathcal{F}$.
For the setting with unknown distributions $\mathcal{D}_{i}$, we use our online framework with binary feedback: in each period, the seller chooses a policy and only observes which customers accepted their offers. As an application of Theorem 4.2, we obtain $\alpha$-regret of $O\left(n\sqrt{kT\log(knT)}\right)$, where $\alpha$ is the best approximation ratio for the SPM instance with known distributions. Below, we list some concrete SPM applications with their best approximation ratios $\alpha$.
  • When $\mathcal{F}=\{S\subseteq[n]:|S|\leq\kappa\}$ (i.e., a uniform matroid of rank $\kappa$), the optimal policy for FSPM can be computed efficiently using dynamic programming; so $\alpha=1$ here. For ASPM under this feasibility constraint, Chakraborty et al. (2010) provide a polynomial-time approximation scheme (PTAS), so $\alpha=1-\epsilon$ for any fixed $\epsilon>0$.
  • When $\mathcal{F}$ corresponds to the independent sets of a general matroid, Yan (2011) provides a $(1-\frac{1}{e})$- approximation algorithm for ASPM.
  • When $\mathcal{F}$ is an intersection of a constant number of matroid, knapsack and matching constraints, Feldman et al. (2021) provide a framework to achieve an $\alpha=\Omega(1)$-approximation algorithm for ASPM.
Previous work by Singla and Wang (2024) obtained online algorithms for FSPM under a uniform matroid of rank one, i.e., the seller can serve only one customer and the customers arrive in a fixed order. Notably, they considered the bandit feedback model, which, in this specific setting, coincides with the binary feedback model. [^5] Although the bandit feedback model only reveals the final revenue, one can recover the identity of the customer who accepted the service by introducing arbitrarily small perturbations to the posted prices. Our result for this case is a $1$-regret of $O\left(n\sqrt{kT\log(knT)}\right)$, which improves upon the $O(n^{2.5}k\sqrt{T}\log T+n^{5}k\log^{2}T)$ bound from  Singla and Wang (2024). We note that this paper also obtained $poly(n)\cdot\sqrt{T}$ regret for continuous distributions (under some regularity assumption).
Note that any SPM policy chooses a threshold (the price $p_{i}$) for each customer $i$. Moreover, the increase in the objective value for any customer $i$ is determined by the binary observation $\mathbf{I}_{\mathsf{X}_{i}\geq p_{i}}$. So, SPM satisfies Assumption 4.1. We now prove monotonicity of SPM. Without loss of generality, for any decision tree (policy) $\Pi$, we assume that the left branch out of a node $(i,p_{i})$ corresponds to customer $i$ accepting while the right branch corresponds to rejection.
Lemma 5.8.
[…]
Proof.
Consider two product distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ such that $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$. We will show that $\mathtt{OPT}(\mathbf{{D}})\leq\mathtt{OPT}(\mathbf{{E}})$ by induction on the number of customers $n$.
Base case ($n=1$). Fix two distributions $\mathcal{D}_{1}\preceq_{\mathsf{SD}}\mathcal{E}_{1}$. Let $p$ be the price offered to this customer under the optimal policy for $\mathcal{D}_{1}$. Then the optimal revenue satisfies:
$\mathtt{OPT}(\mathbf{{D}})=p\cdot\Pr_{\mathsf{X}_{1}\sim\mathcal{D}_{1}}(\mathsf{X}_{1}\geq p)\leq p\cdot\Pr_{\mathsf{X}_{1}\sim\mathcal{E}_{1}}(\mathsf{X}_{1}\geq p)\leq\mathtt{OPT}(\mathbf{{E}}),$
where the first inequality follows from stochastic dominance.
[…]
$\mathtt{OPT}(\mathbf{{D}})$ $=\Pr_{\mathcal{D}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p+V_{L}(\mathbf{{D}}))+\Pr_{\mathcal{D}_{r}}(\mathsf{X}_{r}<p)\cdot V_{R}(\mathbf{{D}})$
$=\Pr_{\mathcal{D}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p-V_{R}(\mathbf{{D}})+V_{L}(\mathbf{{D}}))+V_{R}(\mathbf{{D}}).$ (40)
We now claim that without loss of generality
$p\geq V_{R}(\mathbf{{D}})-V_{L}(\mathbf{{D}}).$ (41)
Indeed, if $p<V_{R}(\mathbf{{D}})-V_{L}(\mathbf{{D}})$ then raising the price to $\bar{p}=V_{R}(\mathbf{{D}})-V_{L}(\mathbf{{D}})$ can only increase the expression (40). Using this, we have:
$\mathtt{OPT}(\mathbf{{D}})$ $=\Pr_{\mathsf{X}_{r}\sim\mathcal{D}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p-V_{R}(\mathbf{{D}})+V_{L}(\mathbf{{D}}))+V_{R}(\mathbf{{D}})$
$\leq\Pr_{\mathsf{X}_{r}\sim\mathcal{E}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p-V_{R}(\mathbf{{D}})+V_{L}(\mathbf{{D}}))+V_{R}(\mathbf{{D}})$
$=\Pr_{\mathsf{X}_{r}\sim\mathcal{E}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p+V_{L}(\mathbf{{D}}))+\Pr_{\mathsf{X}_{r}\sim\mathcal{E}_{r}}(\mathsf{X}_{r}<p)\cdot V_{R}(\mathbf{{D}})$
$\leq\Pr_{\mathsf{X}_{r}\sim\mathcal{E}_{r}}(\mathsf{X}_{r}\geq p)\cdot(p+V_{L}(\mathbf{{E}}))+\Pr_{\mathsf{X}_{r}\sim\mathcal{E}_{r}}(\mathsf{X}_{r}<p)\cdot V_{R}(\mathbf{{E}})\leq\mathtt{OPT}(\mathbf{{E}}),$
Above, the first inequality follows from stochastic dominance and (41). The second inequality follows from the induction. This completes the inductive step. ∎
Using an identical proof as above, we can also prove that FSPM (the fixed order variant) is down-monotone.
@@ Appendix A Sampling-based Algorithms
A stronger monotonicity condition is that of strong monotonicity (Guo et al., 2021), defined next.
@@
A stochastic problem is strongly up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{{E}}$, we have $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{D}}},\mathbf{{x}})]\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$, where $\sigma_{\mathbf{{D}}}$ is the optimal policy for instance $\mathcal{I}$ under distribution $\mathbf{{D}}$. Strong down-monotonicity is defined similarly.
We note that strong monotonicity implies monotonicity. In prior work, Guo et al. (2021) gave optimal sample complexity bounds for stochastic optimization problems that exhibit the strong monotonicity condition. They also proved this property for problems including Prophet inequality and Pandora’s box. We make use of this property in §5, when we apply our result to these problems.
Below, we assume that $f_{max}=1$ by scaling.
Theorem A.1 (Theorem 17 Guo et al. (2021)).
@@
Recall that, in the full-feedback model, we get one sample from each r.v. in every time-step. We now describe the algorithm. The algorithm is straightforward: for each $t=1,2,\ldots T$, we use estimates from the prior t1t-1 time steps to obtain a policy to use for the $t^{\text{th}}$ time-step.
@@
$\mathbb{E}[R_{T}]\leq\mathbb{E}[R_{T}\mid G]\cdot\Pr(G)+\mathbb{E}[R_{T}\mid\overline{G}]\cdot\Pr(\overline{G})\leq O\left(\sqrt{nT\log(nT)}\right)+T\cdot\frac{1}{T}=O\left(\sqrt{nT\log(nT)}\right).$
@@
A significant challenge in the semi-bandit feedback model arises from our lack of control over the r.v.s from which we get samples. One potential strategy to address this issue is to artificially generate samples for an item $i$ by probing item $i$ first in the algorithm’s policy for that period. However, this approach comes with an inherent drawback – probing item $i$ first may result in a poor policy, and so we suffer a high regret in such periods. The standard explore-then-exploit algorithm that first gets $T^{2/3}$ samples for each r.v., and then plays the optimal policy (for the empirical distribution) for the remaining time steps only achieves regret $\widetilde{O}(T^{2/3})$. We are not aware of any generic approach that reduces sample-complexity bounds to regret minimization in the semi-bandit have $o(T^{2/3})$ regret.
@@
We note that Guo et al. (2021) also give (slightly worse) sample complexity bounds for a broader class of problems that need not satisfy any monotonicity property, but have a finite support-size $k$. Using the approaches described above, we can convert these sample-complexity guarantees to obtain $O\left(\sqrt{nkT\log(nT)}\right)$ regret under full-feedback and $O\left((nkT)^{2/3}(\log T)^{1/3}\right)$ regret under semi-bandit (and bandit) feedback.
Appendix B Lower Bounds
We now state some regret lower bound results for monotone problems under the semi-bandit and binary feedback models. We highlight particular applications for which lower bounds are already known, and observe that since these problems can be captured by our framework, the corresponding lower bounds also apply to our setting.
Theorem B.1.
Any semi-bandit learning algorithm for stochastic monotone problems has regret at least $\Omega\left(\sqrt{nT}\right).$
Proof.
We observe that the classical multi-armed bandit (MAB) problem with Bernoulli rewards is a special case of our setting. Indeed, consider the stochastic problem with $n$ independent Bernoulli r.v.s $\{\mathsf{X}_{i}\}_{i=1}^{n}$ where a policy is allowed to probe only one random variable and receives reward $\mathsf{X}_{i}$ for probing $i$. This problem is clearly down-monotone, and it is equivalent to MAB. As shown in Lattimore and Szepesvári (2020), any algorithm for MAB has regret at least $\Omega(\sqrt{nT})$. ∎
Hence, our upper bounds match this lower bound up to a factor of $\sqrt{n}$ and logarithmic terms. The next result, also observed in Singla and Wang (2024), shows that stochastic problems under binary feedback with continuous distributions are fundamentally harder to learn than with discrete distributions.
Theorem B.2.
Any learning algorithm for stochastic monotone problems with binary feedback and continuous distributions has regret at least $\Omega\left(T^{\frac{2}{3}}\right).$
Proof.
Consider the fixed-order SPM under a uniform matroid of rank one, which is a down-monotone stochastic problem. Leme et al. (2023) proved an $\Omega(1/\epsilon^{3})$ query complexity lower bound for this problem under continuous distributions with binary-threshold queries. This directly implies an $\Omega(T^{2/3})$ regret lower bound for online learning for SPM with binary feedback. Suppose (for a contradiction) that there is an online algorithm with regret $o(T^{\frac{2}{3}})$. Then after $T$ rounds, we can achieve $o(T^{-\frac{1}{3}})$ per-round regret, which means that we can obtain a policy with $\epsilon$ additive error in $T=o(1/\epsilon^{3})$ rounds, which contradicts the sample complexity result. ∎
Acknowledgement
We thank Ziyun Chen for pointing out some missing assumptions in our result for continuous distributions, which was stated in an earlier version of this paper. This observation lead us to the corrected statement and proof of Theorem 3.7.